OpenStudy (anonymous):

y=4x2+28x+47 se completing the square to find the x−intercepts and the vertex of the graph. Answer as ordered pairs.

OpenStudy (anonymous):

you do not need to complete the square to find the \(x\) coordinate of the vertex, it is much easier to use \(x=-\frac{b}{2a}\) which in your case is \(-\frac{28}{2\times 2}\)

OpenStudy (anonymous):

sorry i misread it, it is \[-\frac{28}{2\times 4}\]

OpenStudy (anonymous):

HOW DID U GET THAT AND IS Y THAT THE FINAL ANSWER